# Rapid Muon Tomography for Border Security

### v1

@@ III Results
Illegal trafficking of nuclear materials poses a significant challenge for terrorism prevention. We have modeled a scenario in which high-Z materials are hidden inside steel pipes, which are standard customs goods but can serve as effective shielding to conceal nuclear weapons materials.
@@
Figure 9: (a) Shipping container GEANT4 model with SNM cubes (red). (b) Reconstructed 3 min PoCA image of SNM cubes hidden in steel pipes. Noise from steel tubes removed using adaptive thresholding algorithm.
Figure 10: XY and XZ projections of reconstructed image for the scenario of three high-Z material cubes hidden inside steel pipes: (a),(b) for 1 minute scanning time and (c),(d) for 3 minute scanning time.
@@
In this scenario, a shipping container is loaded with 168 steel pipes, each with a diameter of 16 cm, length of 450 cm and a thickness of 1 cm with a total weight of $\approx$27 tons. Reconstructed image for the scenario of three high-Z material cubes hidden inside steel pipes. Concealed within the pipes are three 10 cm3 Special Nuclear Material (SNM) cubes (marked in red, see Figure 8), strategically hidden within the dense steel structure to evade detection.
The ’Prompt’ cargo inspection algorithm, applied over a 10-second period, was used to detect the presence of high-Z material in the cargo. As shown in Figure 9 (a), for the 100,000 muons simulated, the reconstructed distribution - obtained using the PoCA method - includes 5,000 data samples for a shipping container with steel pipes (blue data points) and 5,000 data samples for steel pipes containing three hidden 10 cm3 SNM cubes (red data points). As observed, both distributions overlap.
To improve detection, a longer scanning time was evaluated for identifying Special Nuclear Material (SNM) concealed within dense steel pipes using an adaptive thresholding algorithm. This method effectively filters out noise from the steel pipes, enhancing contrast between high-Z materials and the surrounding cargo structure, leading to more precise SNM identification. Figure 9(b) presents the reconstructed 3D image. Figure 10(a) shows the reconstructed 1-minute tomographic image, generated from 1 million muons using the CRY muon generator. While the adaptive thresholding algorithm has significantly reduced noise from the steel tubes, some residual noise remains. The reconstructed tomographic image for a 3-minute scanning time, shown in XY and XZ 2D projections (Figure 10(b)), provides a clearer view of the SNM cubes hidden within the steel pipes. The noise from the steel tubes has been effectively removed, enhancing the visualization and detection of SNM.
IV Discussion
The investigated two-stage approach for cargo inspection across different scenarios can be considered a viable method for implementation in customs practices, enhancing the efficiency and accuracy of cargo screening. In the first stage, a scanning time of 10 to 20 seconds is used to analyze the cargo’s scattering and absorption rates. In the second stage, a 3D image is reconstructed within 1 to 5 minutes, enabling the detection of contraband materials through post-processing techniques. In the scenarios considered, branded textile products substituted for declared goods were detected within 10 seconds using the prompt approach, which combines scattering and absorption data, and visualized within 30 seconds. Hidden explosives were detected within 10 seconds using the same prompt approach, with visualization achievable within 5 minutes. The SNM concealed within dense cargo, specifically steel pipes, could not be identified using the prompt approach. However, detection and visualization of SNM were achieved within 3 minutes using an adaptive thresholding algorithm that effectively filters out noise from the steel pipes.
The scenarios modeled in this study represent only a subset of real-world smuggling cases. However, they address key categories of contraband, including low-density and low-atomic-number materials used to disguise the misdeclaration of high-value goods, a common tactic in fraudulent trade practices. Hazardous and volatile materials, which pose significant risks to public safety and require rapid, accurate identification. Nuclear material smuggling, a major concern in non-proliferation efforts, where high-Z materials are deliberately concealed within dense cargo structures. Future work should expand the range of simulated scenarios to include more complex concealment methods, such as layered shielding techniques, mixed-material cargo, and varying cargo densities.
Future work should expand the range of simulated scenarios to incorporate more complex concealment methods, such as layered shielding techniques, mixed-material cargo, and varying cargo densities. In practical applications, the large volume of trade facilitates the creation of an extensive database of 3D cargo images. This enables the integration of artificial intelligence (AI) to train models on real cargo images for accurate image categorization, allowing for rapid detection of discrepancies between declared cargo and measured data, reducing response times and improving inspection efficiency.
V Conclusion
@@
This study demonstrates the effectiveness of a two-stage cargo screening system for real-time cargo composition estimation. Our integrated scattering and absorption approach has proven to be a highly sensitive method for cargo analysis, offering significant potential for security applications.
Applying this method to multiple Monte Carlo-simulated scenarios highlights its ability to rapidly detect discrepancies between declared cargo contents and measured scattering and absorption rates, enabling customs officers to respond swiftly. Additionally, advanced 3D image processing techniques, including median and Gaussian filtering and adaptive thresholding, enhance image clarity by reducing noise, identifying anomalies, and improving the detection of concealed hazardous materials.
The highly sensitive data processing algorithms developed in this study establish muon tomography as a powerful tool that complements X-ray scanning procedures, providing a non-invasive, high-precision solution for border security and cargo inspection.
@@ Acknowledgements.
This work was funded by the EU Horizon 2020 Research and Innovation Programme under grant agreement no. 101021812 (“SilentBorder”).

### v2

@@ III Results
Refer to captionRefer to caption
Figure 9: (a) 3D visualization of three detected SNM cubes using the DBSCAN algorithm. The cubes are marked with different colors corresponding to the identified clusters. XY, YZ and XZ projections of tomographic image for 1 minute scanning time.
@@
Illicit trafficking of nuclear materials remains a critical security concern. We modeled a scenario where high-Z Special Nuclear Material (SNM) is concealed within steel pipes, common cargo items that can act as effective shielding. The GEANT4 model of container with pipes was created. We simulated one million muons sampled on the surface 10 $\times$ 10 cm2 using CRY particle generator, which corresponding to a one-minute scan, to evaluate the system’s detection capability. In this scenario, a shipping container is loaded with 166 steel pipes, each with a diameter of 16 cm, length of 500 cm and a thickness of 1 cm. Concealed within the pipes are three 10 $\times$ 10 $\times$ 10 cm3 Special Nuclear Material (SNM) cubes (marked in red in Figure 8), strategically hidden within the dense steel structure to evade detection.
The simulated data were processed using PoCA imaging algorithm. Performing image processing an adaptive thresholding algorithm was employed to suppress noise originating from the surrounding steel structures. This significantly improved image clarity and reduced false positives. Subsequently, the DBSCAN clustering algorithm was applied to automatically identify and localize the three SNM objects within the reconstructed volume. Figure 9(a) presents the reconstructed 3D image processed using the DBSCAN algorithm. Three SNM cubes are detected and marked with different colors corresponding to the identified clusters. Figures 9(b)–(d) show the XY, YZ, and XZ 2D projections, respectively. These views provide a clearer visualization of the SNM cubes hidden within the steel pipes. Noise from the surrounding steel structure has been effectively suppressed, enhancing both the image clarity and the detectability of SNM.
These results demonstrate the strong potential of real-time muon tomography combined with intelligent post-processing for the rapid detection of nuclear threats in dense and shielded cargo environments.
IV Discussion and Conclusions
@@
This study demonstrates the effectiveness of a two-stage cargo screening system based on muon tomography for real-time cargo composition estimation. The integration of scattering and absorption measurements offers a highly sensitive, non-invasive method for analyzing a wide range of cargo types, with strong potential for enhancing customs screening procedures.
The proposed approach involves a rapid initial scan (10–20 seconds) to evaluate scattering and absorption rates, followed by a second stage in which a 3D image is reconstructed within 1 to 5 minutes. Post-processing techniques are then applied to detect and localize concealed contraband materials with high precision.
Monte Carlo-simulated scenarios confirm the system’s ability to rapidly detect discrepancies between declared cargo content and measured physical properties. Categories of detected contraband include:
  • Low-density, low-Z materials, commonly used to disguise the misdeclaration of high-value goods.
  • Hazardous and volatile substances, which pose immediate safety risks and require prompt identification.
  • Special Nuclear Materials (SNM), often deliberately shielded within dense cargo structures to evade conventional detection—posing a critical non-proliferation concern.
Advanced 3D image processing techniques, including median and Gaussian filtering and adaptive thresholding, significantly improve image clarity by suppressing structural noise and enhancing anomaly detection. These methods have proven effective in revealing both low-Z and high-Z concealed threats, even when embedded in complex or dense cargo configurations.
The highly sensitive data processing algorithms developed in this work establish muon tomography as a powerful complement to traditional X-ray imaging systems. Especially in cases involving shielded or irregular cargo where X-rays may fail, muon-based systems offer a robust, high-accuracy, and non-destructive solution for modern border security and cargo inspection.
Future work should focus on integrating machine learning for anomaly detection and expanding real-world testing to further validate performance under operational conditions.
@@ Acknowledgements.
This work was partially supported by the EU Horizon 2020 Research and Innovation Programme under grant agreement no. 101021812 (“SilentBorder”).